OpenStudy (anonymous):

I need help with this following problem, the question is: Solve Using Transformations: 7(8 - 3) + x = 42. What does this mean? It's on my study guide for a test please help!

OpenStudy (amistre64):

it means that you subtract 7(8-3) from both sides and solve for x: x = 42 - 7(8-3) x = 42 - 7(5) x = 42 - 35 x = 7

OpenStudy (anonymous):

Thank you sir =].

OpenStudy (amistre64):

yw :)